Operations Report – January 24th

SOL: 3
Name of person filing report: Rob Hunt + Scott Dorrington
Non-nominal systems: None
Notes on non-nominal systems: None
Spirit rover used: Scott EVA #5
Hours: 242.4 (start) – 242.6 (end)
Beginning charge: 100%
Ending charge: 66%
Currently charging: No
Opportunity rover used: Alex and Claire EVA #5
Hours: 156.5 (start) – 156.8 (end)
Beginning charge: 100%
Ending charge: 69%
Currently charging: No
Curiosity rover used: N/A not used
Beginning charge:
Ending charge:
Currently charging:
Perseverance rover used: N/A not used
Beginning charge:
Ending charge:
Currently charging:
General notes on rovers:
Opportunity – nominal
Spirit – nominal
Summary of Hab operations:
WATER USE: 427.5 gal.
Water (static tank): 393.2 gal.
Static tank pipe heater (on or off): On
Static tank heater (On or off): On
Toilet tank emptied: no
Summary of internet: Nominal
Summary of suits and radios:
Donning area negligibly re-arranged for better circulation space during donning and doffing,
Calico ‘head’ covers for suits vary in circumference so they were sorted and number-matched to ensure best fit,
Battery voltage on all suits is now recorded before and after EVAs. This has revealed at least two that lose approx 2V during an EVA. i.e. approx 13.5V to 11.5V.
Two-piece suits
Suits #1 – #6 were checked for structural integrity. Some minor repairs made e.g. loose screws tightened (not overtightened)
Suits #1 – #6 chargers labelled/placed according to suit numbers,
Charger with label #3 missing and charger with label #12 was relabelled.
One-piece suits
Suit #10 – strap-ends tidied with zip-ties to make donning less confusing. This is planned for the other suits,
Suit #11 (small size) – power issue resolved. Problem was the toggle switch which has a ‘neutral’ setting between ‘Off’ and ‘On’. It is the only one of the 11 suits with this type of switch. Best solution is to swap out with same type switch as others. Other solutions might be general team awareness and/or a check list for donning area.
Summary of GreenHab operations: put notes here
WATER USE: 6.5 gallons
Heater: Off (broken). Space heater on overnight
Supplemental light: On
Harvest: 14g basil, 21g dill, 38g kale, 57g red cabbage,
Summary of ScienceDome operations: put notes here
Dual split: nominal
Summary of RAM operations:
Collated and photographed existing stock of ‘tape’,
Collated and photographed existing stock of cable ties,
Used 2 x 3” screws to repair door striker on Room 4
Summary of any observatory issues: Observatory not used.
Summary of health and safety issues:
Hazards identified:
Questions, concerns and requests to Mission Support:
Rolls of fabric tape, insulation tape, clear packing tape please (see pic of current stock),
Large cable ties please (for the walkway tarp)
Bread knife
2L plastic drink jug
Silicon baking spatula

Copyright © The Mars Society. All rights reserved. | Main Site